Norma invested the start-up capital for her Internet business in two hedge funds. After one year one of the funds returned 5% and the other returned 6% for a total return of $5880. If the amount on which she made 6% was $10,000 larger than the amount on which she made 5%, then what was the original amount of her start-up capital?
Answer by Fermat(127)

Let A be the amount invested at 5%.
Let B be the amount invested at 6%.
return on A is 0.05A
return on B is 0.06B
total return is $5,880
.: 0.05A + 0.06B = 5880 -----------------(1)
B is greater than A by $10,000, i,e.
B = A + 10000
0.06B = 0.06A + 600 ---------------------(2)
substituting for B, from (2), into (1),
0.05A + 0.06A + 600 = 5880
0.11A = 5280
A = 48000
B = 58000
Total Startup fund = A + B = $106,000
======================================